AO Commissioning on 13 June 2011

Observer(s): D. Miller, JC. Guerra, K. Boutsia, E. Pinna, L Busoni, A. Puglisi, S. Esposito, J. Hill, F. Quiros-Pacheo, J. Argomedo (LBTO)

Telescope Operator: S. Allanson (LBTO)

SW Support: N. Cushing (Tucson)

Telescope Support: M. Midkiff (LBTO)

Instrument Support: T. Shih(LBTO)

Summary

Details

02:30 twilight flats J, H, Ks (10sec)

02:40 pointing correction and them going for a standard

02:49 iif_preset, FS23, MODE=ACQUIRE, GS=0, ROTMODE=IDLE

02:53 iif_preset, FS23, MODE=ACQUIRE, TS=8, GS=8, ROTMODE=IDLE
  • too bright to see the standard we try to go to a nearby brighter star

02:55 iif_preset, FS23, MODE=ACTIVE, TS=0, GS=8, PA=-110
  • Ks, J, H,
  • iif_offset, 6,-1, COORD='DETXY'
  • iif_offset, 6,-11, COORD='DETXY'

03:29 iif_preset, FS23, MODE=ACTIVE, TS=0, GS=12, PA=-110
  • iif_offset, -6,-11, COORD='DETXY'
  • iif_offset, -5,-1, COORD='DETXY'

03:43 iif_preset, Pisces_13, MODE=ACQUIRE, GS=0, ROTMODE=IDLE

03:49 iif_preset, Pisces_13, MODE=ACE, GS=3, RS=0, PA=60
  • the AO system is stopped in its configuration stage
 
===> Error in irc_presetTelescope
        err = Resmsg[0]  = PresetTelescope result status: Error
        err = Resmsg[1]  = Stop[90] complete
        err = Resmsg[2]  = PresetAO[91] failed - Command timeout - -13107: Arbitrator interface generic error

03:55 iif_preset, Pisces_13, MODE=ACE, GS=3, RS=0, PA=60
  • DIMM 0.9''
  • same problem as before!

03:59 iif_preset, Pisces_13, MODE=ACE, GS=3, RS=0, PA=60
  • irc_runao
  • AOS timeout error - AO system still trying to close the loop

04:08 iif_preset, Pisces_13, MODE=ACE, GS=3, RS=0, PA=60
  • preset failed again - WFS susystem is not available

04:17 iif_preset, Pisces_13, MODE=ACE, GS=3, RS=0, PA=60
  • this time we have a TCS error - because they turned the mirror power off and this also turned the hexapod off
  • hexapod had to be initialized and homed by T.O.

04:20 iif_preset, Pisces_13, MODE=ACE, GS=3, RS=0, PA=60
  • presetAO failed

04:27 iif_preset, Pisces_13, MODE=ACE, GS=3, RS=0, PA=60
  • irc_runao - SUCCESS!!!!!!! finally!

06:19 iif_offset, 0, -5, COORD='DETXY' Failed, loop gain was not set to zero and mirror followed offset
  • Alfio does some work

06:30 iif_preset, Pisces_13, MODE=ACE, GS=3, RS=0, PA=60
  • no irc_runao sent because a "new" problem wth AO system. What?

06:40 iif_preset, Pisces_13, MODE=ACE, GS=3, RS=0, PA=60
  • irc_runao success
  • closed loop immediately failed. Tip/Tilt mirror was not syncronized with (?) so caused the loop to open. Next preset and acquire were successful with no changes to AO system/TT mirror.

06:45 iif_preset, Pisces_13, MODE=ACE, GS=3, RS=0, PA=60
  • irc_runao Success
  • iif_offset, 0, -5, COORD='DETXY'
  • iif_offset, 0, -40, COORD='DETXY' Failed. Light not on Pyramid. See ref star is ~2" from hotspot on TV. Might be a scale or rotation problem.

06:55 iif_preset, Pisces_9, MODE=ACQUIRE, GS=0, ROTMODE=IDLE Success

07:05 iif_preset, Pisces_10, MODE=ACE, GS=7, RS=0, PA=110
  • the wrong filed - the previous time they used this to take dark frames

07:09 iif_preset, Pisces_9, MODE=ACE, GS=19, RS=0, PA=110

07:12 iif_preset, Pisces_9, MODE=ACE, GS=19, RS=0, PA=110
  • send preset again because the guide star is not on the hot spot
  • now it is centered!
  • irc_runao - it send an timout error on the idl frame but the loop is actually closed and they go on with the observations
  • we tried to do a rotational offset from iraf and AOS complained and then it failed because it was off-source too long

07:24 iif_preset, Pisces_9, MODE=ACE, GS=19, RS=0, PA=90
  • the preset failed - the shell is ripped

07:28 iif_preset, Pisces_9, MODE=ACE, GS=19, RS=0, PA=90
  • irc_runao
  • they are trying to have the same reference star as the previous time.

07:59 loop opened: firmware problem on BCU on the WFS

08:27 observe, 'test.org', logname, retval First test of observing script Success, with only preset part
  • irc_runao

test.obs = "preset, OBJECT=Pisces_10, MODE=ace, GS=11, RS=0, PA=90"

This sends a preset in ACQUIRE and IDLE mode

08:42 loop opened and shell ripped - unknown reason

08:50 observe, filein, logname, retval - successful
  • runao - failed because of AOS timeout but the loop closed
  • preset parameters:
          # Target star: 0   Pisces_10
            17:17:12.8030 +43:05:41.960 RADEC 2000.0 J2000 0.0 0.0 10.50 R 0.00 B_V 1.65
          # Guide star: 11
            17:17:33.8179 +43:04:41.887 RADEC 2000.0 J2000 0.0 0.0 12.43 R 0.98 B_V 0.63
          # AO Reference star: 0
            17:17:12.8030 +43:05:41.960 RADEC 2000.0 J2000 0.0 0.0 10.50 U 0.00 B_V 0.5
      => PresetTelescope, PA=90.0000, ROTMODE=POSITION, MODE=ADAPTIVEACE
    

09:25 iif_preset, Pisces_9, MODE=ACE, GS=19, RS=0, PA=90
  • irc_runao - failed because there were too many stars and could not choose a reference star
  • irc_runao - failed again for the same reason
  • DIMM 1.0 - 1.1''
  • irc_runao - error because of AOS timeout but the loop is closed

10:21 iif_preset, Wref11new, MODE=ACQUIRE, GS=0, ROTMODE=IDLE

10:31 iif_preset, Wref11new, MODE=ACE, TS=7, GS=1, RS=7, PA=-90
  • irc_runao

11:03 iif_preset, Wref11new, MODE=ACE, TS=0, GS=6, RS=0, PA=-90
  • irc_runao - runao did not give any error BUT the loop opened (probably due to a problem with the gain) and the mirror ripped

11:13 iif_preset, Wref11new, MODE=ACE, TS=0, GS=6, RS=0, PA=-90
  • irc_runao - again shell ripped without giving any error to the iif
  • DIMM - 1.2-1.3''

11:20 iif_preset, Wref11new, MODE=ACE, TS=0, GS=6, RS=0, PA=-90
  • irc_runao - this time it took some time but it was successful
  • DIMM - 1.5''

11:40 iif_offset, 0, 0 , COORD='DETXY', ROT=5
  • it did not give an error but no rotation took place - we repeat
  • iif_offset, 0, 0 , COORD='DETXY', ROT=5
  • again nothing seems to rotate

11:45 we rotate 5 degrees using iraf LBT tools - this time the offset is successful and the loop did not open
  • we repeat and we give 5 more degrees - the rotation was successful again - we keep rotating 5 degrees
  • check iif_offset routine to see why ROT=? does not send the rotation

11:57 - preset failed - the offset completed before the slew was done - End of science!

11:58 iif_preset, WT10_304, MODE=ACQUIRE, GS=0, ROTMODE=IDLE - pointing for flats
  • sky flats in J,H,Ks,FeII,H2,Brg

12:27 - End of night

Morning twilight. Sky flats: H,Ks,J,FeII,H2,Br_gm, Darks(1,2,3,5,10,15,30,60sec) images filter. .0687 H 10sec 4 frames .0693 H 10 frames 3 frames .0697 ks 4frame .0700 J 10sec 1 frames .0705 J 4sec 3 frames. .0712 ks 7sec 3 frames. .0713 FeII 7sec 1frame .0714 FeII 30sec 3 frames .0717 H2 10 sec 3 frames. .0720 H2 30 sec 3 frames. .0723 Br_gm 10 sec 4 frame .0727 Br_gm 30sec 3frames.

summary standard and flats
i files con .hdr.fits can be found in the directory /OldRepository/PISCES_Data/20110613/ and most of the frames until .0300.fits have no .hdr.fits and can be found in /newdata/warning

FS23 various offsets filter
pisces.20110613.0048.hdr.fits - *.0051.hdr.fits Ks
pisces.20110613.0052.hdr.fits - *.0056.hdr.fits H
pisces.20110613.0041.fits - *.0042.fits Ks
pisces.20110613.0043.fits - *.0047.fits J
pisces.20110613.0057.fits - *.0062.fits H
pisces.20110613.0063.fits - *.0066.fits Ks
pisces.20110613.0067.fits - *.0075.fits J
pisces.20110613.0076.fits - *.0078.fits Ks
pisces.20110613.0079.fits - *.0085.fits H
pisces.20110613.0086.fits - *.0088.fits Ks
pisces.20110613.0089.fits - *.0094.fits J
pisces.20110613.0095.fits - *.0097.fits Ks
pisces.20110613.0098.fits - *.0100.fits H

evening flats - lowish counts filter
pisces.20110613.0005.fits - *.0010.fits Ks
pisces.20110613.0012.fits - *.0026.fits J
pisces.20110613.0027.fits - *.0037.fits H

dawn flats - low and high counts filter
pisces.20110613.0686.hdr.fits - *.0695.fits H
pisces.20110613.0696.hdr.fits - *.0699.fits Ks
pisces.20110613.0700.hdr.fits - *.0707.fits J
pisces.20110613.0708.hdr.fits - *.0714.fits Ks
pisces.20110613.0715.hdr.fits - *.0718.fits FeII 1.64 um
pisces.20110613.0719.hdr.fits - *.0722.fits H2 2.122 um\xA0
pisces.20110613.0723.hdr.fits - *.0729.fits Br-g 2.166 um

dark frames
dark frames exptime
pisces.20110613.0730.hdr.fits - *.0734.fits 1sec
pisces.20110613.0735.hdr.fits - *.0739.fits 3sec
pisces.20110613.0740.hdr.fits - *.0744.fits 5sec
pisces.20110613.0745.hdr.fits - *.0749.fits 2sec
pisces.20110613.0750.hdr.fits - *.0754.fits 8sec
pisces.20110613.0755.hdr.fits - *.0759.fits 10sec
pisces.20110613.0760.hdr.fits - *.0764.fits 15sec
pisces.20110613.0765.hdr.fits - *.0769.fits 30sec

-- KonstantinaBoutsia - 13 Jun 2011
Topic revision: r10 - 10 Feb 2012, MichaelWagner
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback